# Enhanced indexation using both equity assets and index options

### v1

@@ 1 Introduction
Enhanced indexation refers to the process of assembling a financial portfolio in an attempt to (hopefully) outperform a specified marked index. Major market indices here include the S&P 500 in the USA and the FTSE 100 in the UK.
A common approach to formulating the problem of deciding an enhanced indexation portfolio is a historic lookback approach. To look into known asset prices (and returns) in the immediate past and ask the question what would have been the best set of assets to hold over this past period, and in what proportions, so as to best outperform the index”. If we can answer this question (e.g. using some optimisation model) then it seems reasonable to hold that portfolio into the immediate future, especially since this removes the need to try and forecast future prices.
For example if over the immediate past putting 20% of our wealth into Amazon stock contributes to the best portfolio to outperform the S&P 500 in the immediate past then it seems reasonable to put 20% of our wealth into that stock to hold into the immediate future.
@@
An implicit assumption here is that any asset considered for inclusion in the enhanced indexation portfolio based on the immediate past will also exist into the immediate future. For assets in the major indices such as the S&P 500 this assumption is valid. However this assumption is no longer true if we consider options.
@@
Briefly here (since we discuss options in greater detail below) financial options exist relating to the market index value. For example a call option relating to the S&P 500 would have a specified index value ($I$, say), at a specified time in the future (so the option has a specified life before expiry). If we choose to pay to buy this option now then we can, if we wish, sell it over its life. At the end of its life (assuming we have not sold the option in the meantime) if the market index is above $I$ then we receive a cash amount depending upon the difference between the market index value and $I$. If the market index value is below $I$ then we receive nothing. So here the return on the option can be calculated using the price paid and the sum received (either on sale, or at end of life).
@@
But because an option has a specified life, and indeed can be sold over its life, the assumption referred to above, namely any asset considered for inclusion in the enhanced indexation portfolio based on the immediate past will also exist into the immediate future, is no longer true.
For example, suppose we know (somehow) that a good enhanced indexation portfolio in the immediate past would have put 20% of our wealth into a S&P 500 call option four weeks ago, and sold it after a week, so three weeks ago. Does this help us in knowing what to do in the immediate future? Clearly not.
@@
The difficulty here is that equity assets have a “long life”, options have a “short life”. In this paper we propose an approach based on option strategies that enables us to overcome this difficulty, and hence consider both equity assets and options for inclusion in an enhanced indexation portfolio. The simple justification as to why we might consider index options in enhanced indexation is that index options on major indices are widely available. As such including them in a portfolio seeking to outperform an index can (depending upon the decisions taken) contribute to achieving better outperformance than using equity assets alone.
@@
An option strategy for a known set of options is a specified set of rules which detail how these options are to be traded (bought, rolled over, sold) depending upon market conditions. Applying an option strategy to the immediate past means that we can evaluate it, so effectively it is now an artificial asset with known performance. Because it is a strategy it has a “long life”, so can be chosen for inclusion in the enhanced indexation portfolio and applied in the future. Each option strategy contributes one artificial asset. These assets, together with equity assets, can be used in any optimisation approach for enhanced indexation. In this paper we make use of an approach based upon linear programming and second-order stochastic dominance.
@@
The structure of this paper is as follows. In Section 2 we review the relevant literature as to enhanced indexation. In Section 3 we discuss the varying options that are available and how they are valued (priced), with special regard to index options for the S&P 500. We define what we mean by an option strategy, giving an example S&P 500 option strategy and illustrate how we can evaluate such a strategy. We discuss how we can use option strategies in any optimisation based approach to enhanced indexation and state what we believe to be the contribution of the paper.
@@ 2 Literature review
In this section we review the literature relating to enhanced indexation. To do this we made use of Web of Science (http://www.webofscience.com), which is a standard reference database for the refereed research literature. Using this database with the search terms “enhanced indexation” or “enhanced index tracking” yielded over 70 papers. Since the primary purpose of this paper is to focus on the use of equity assets together with options for enhanced indexation we will only discuss here papers that, in our view, are of more significance. We would also mention here the recent work of Silva and de Almeida Filho (2024) who gave a comprehensive literature review relating to index tracking. As part of that review they highlight papers which have extended approaches for index tracking to enhanced indexation.
We would note here that using Web of Science with the search terms “enhanced indexation” and “enhanced index tracking” together with “options” gave no relevant papers. Hence, to the best of our knowledge, this paper is the first in the literature to address enhanced indexation using both equity assets and options.
@@
Early work dealing with enhanced indexation was given by Alexander and Dimitriu (2005) who made use of cointegration and Dose and Cincotti (2005) who clustered stocks based on a distance measure between two stocks defined using the difference between their stock price histories (amongst other approaches). Canakgoz and Beasley (2009) presented a mixed-integer programming formulation that includes transaction costs, a constraint limiting the number of stocks that can be purchased and a limit on the total transaction cost that can be incurred. Their approach makes use of the linear regression between portfolio returns and index returns. Computational results were presented using a standard optimisation package (Cplex).
@@
Li et al. (2011) proposed a bi-objective formulation, minimising a tracking error function related to returns below the index, whilst maximising portfolio return and limiting transaction costs. Computational results were presented using an immunity-based solution approach. Meade and Beasley (2011) presented a paper that makes use of a modified Sortino ratio to produce portfolios of specified cardinality, but with an upper limit on the holding in any particular asset. Computational results were presented using a genetic algorithm.
@@
Guastaroba and Speranza (2012) presented a mixed-integer programming formulation that includes transaction costs, a constraint limiting the number of stocks that can be purchased and a limit on the total transaction cost that can be incurred. Their objective is to minimise the absolute deviation between a scaled increased index value and the portfolio value. Computational results were presented using kernel search and a standard optimisation package (Cplex).
Mezali and Beasley (2013) presented a mixed-integer programming formulation that includes transaction costs, a constraint limiting the number of stocks that can be purchased and a limit on the total transaction cost that can be incurred. Their objective is based on quantile regression. Computational results were presented using a standard optimisation package (Cplex).
@@
Roman et al. (2013) made use of second-order stochastic dominance (SSD), aiming for a portfolio that is second-order stochastic dominant with regard to the index. Their solution procedure involves the use of cutting planes as given by  Fabian et al. (2011a, b) who introduced a cutting plane reformulation of Roman et al. (2006). Computational results were presented using a standard optimisation package (FortMP).
Valle et al. (2014) presented a three-stage mixed-integer programming formulation that includes transaction costs, a constraint limiting the number of stocks that can be purchased and a limit on the total transaction cost that can be incurred. The first two stages relate to a regression of portfolio return against time, whilst the third stage relates to minimising transaction cost. Computational results were presented using a standard optimisation package (Cplex).
@@
Acosta-Gonzalez et al. (2015) presented an approach based on optimising cointegration between the portfolio chosen and the benchmark index. They made use of genetic algorithms. For enhanced index tracking they artificially increase the benchmark index. Bruni et al. (2015) proposed a linear bi-objective optimisation approach that maximises average excess return and minimises the maximum underperformance with respect to the index. Computational results were presented using MATLAB and the TOMLAB/Cplex toolbox.
Filippi et al. (2016) presented a bi-objective mixed-integer programming formulation balancing excess return against the absolute deviation between a scaled index value and the portfolio value. Computational results were presented using kernel search and a standard optimisation package (Cplex). Guastaroba et al. (2016) presented a linear programming formulation based on the Omega ratio due to Keating and Shadwick (2002). This ratio involves setting a predetermined threshold and then partitioning portfolio returns into those below the threshold (losses) and those above the threshold (gains). The Omega ratio is defined as the probability weighted gains divided by the probability weighted losses. Computational results were presented using a standard optimisation package (Cplex).
@@
Clark et al. (2019) presented an approach based upon reweighting the assets in the benchmark index. Their approach makes use of the concept of cumulative utility area ratio, which measures the incremental utility of one asset compared to another, and the analytic hierarchy process Saaty (1980, 1990).
Guastaroba et al. (2020) proposed bi-criteria optimisation models using weighted multiple conditional value-at-risk, defined as a weighted combination of multiple CVaR measures. Computational results were presented using a standard optimisation package (Cplex). Li et al. (2021) presented a nonlinear enhanced indexation model with an explicit objective to track and outperform the naive diversification $(1/N)$ strategy. Computational results were based on simulation.
Xu et al. (2022) make use of group-sparse optimisation, with dynamic parameter adjustment being based on either a hidden Markov model or an artificial neural network. Allen-Zhao et al. (2024) presented an approach based on robust CVaR and group-sparse optimisation. Computational results were presented using standard optimisation packages (Cplex and Gurobi).
Dai and Li (2024) make use of a novel neural network together with a hidden Markov model with Gaussian noise to capture regime switching. Cesarone and Puerto (2025) presented an approach based upon ordered weighted average in conjunction with stochastic dominance. The weighted average is based on the sum of $k$ maximal CVaRs. Computational results were presented using a standard optimisation package (XPRESS).
@@ 3 Options
In this section we discuss the varying options that are available and how they are valued (priced), with special regard to index options for the S&P 500, which we use for our computational work below. We then define what we mean by an option strategy, giving an example S&P 500 option strategy and illustrate how we can evaluate such a strategy. We go on to discuss how we can use option strategies in any optimisation based approach to enhanced indexation and state what we believe to be the contribution of the paper as well as discuss some extensions of the work presented here. Finally we discuss the choice of an optimisation based enhanced indexation approach.
@@ 3.1 Options and option pricing
A call option on an underlying asset gives the holder of the option the right (but not the obligation) to buy the asset at a fixed price, called the strike (or exercise) price. If this right can only be exercised at a specific (expiration) time in the future then it is known as an European call option. If it can be exercised at any time up to expiration then it is known as an American call option.
A put option is the right to sell the asset, otherwise it has the same characteristics as a call option (so a strike price, expiration time, European or American).
@@
Even if an option can only be exercised at expiration it can be bought and sold at any time. So for example if we buy a call or put option now we can sell it at any time between buying it and expiration.
@@
However options also apply if we have intangible assets, such as the S&P 500 index. As we cannot “buy” any of the index an option such as this is cash settled. So the strike price relates to the level of the index. Considering a call option, if at expiration the index is below the strike price the option is worthless. However if at expiration the index is above the strike price then the option is cash settled, so the option holder receives a cash amount linearly related to the difference between the current index level and the strike price.
@@ 3.2 Defining an option strategy
On a technical note here although we use phrases such as “buy the option” we mean buy a number of units of the same option. Since the sums of money involved are likely to be large we can ignore any issues relating to fractional option unit purchase.
@@
In other words an option strategy is equivalent to an “artificial asset” that can be treated just like an equity asset in any current approach for enhanced indexation.
@@ 3.3 An example S&P 500 option strategy
  • In terms of expiration dates the two most liquid S&P 500 options are: Monthly, expiring on the third Friday of every month and EOM, expiring on the last business day of each month.
@@
To illustrate how we rollover suppose the rollover rule above is active. So we are selling the current put option and buying another put option. For the purchased option we need to decide the lifetime and the exercise price. As the rollover rule involves $L<20$ days to invoke sale of the current option then for the purchased option we choose the next available expiration (monthly or EOM) 20 or more days in the future, giving the purchased option a lifetime of expiration time minus current time.
In terms of the exercise price $E$ for the purchased OTM option then as we have a $3\%$ moneyness target and since for a OTM put option $E<F_{t}$ (see Table 1) we have $100(F_{t}–E)/F_{t}=3$ from Equation (6), equivalently $E=0.97F_{t}$. Here $F_{t}$ is calculated from Equation (5) using the lifetime as decided above for the purchased option. As exercise prices are integer multiples of 5 we round $E$ to the nearest multiple of 5.
@@ 3.4 Example S&P 500 option strategy evaluation
In period 2 the 30 day index return is $\leq-5\%$ so we buy the option at 90 so in total $1000u/90$ options are bought, with 21 days to expiration.
@@
In period 4 the options held in period 3 are now 19 days to expiration so we rollover, selling our current holding of $1000u/90$ options for 80. For the new option the number of days to expiration is 29 (e.g. EOM). Since, based on this new lifetime of 29 days, $F_{t}=6379$ the exercise price $E=0.97F_{t}$, which after after rounding to the nearest multiple of 5 is 6190. The new option costs 75 and we hold $(80{\times}1000u)/(90{\times}75)$ units of that option.

### v2

@@ 1 Introduction
Enhanced indexation refers to the process of assembling a financial portfolio in an attempt to outperform a specified marked index. Major market indices here include the S&P 500 in the USA and the FTSE 100 in the UK.
A common approach to addressing the problem of deciding an enhanced indexation portfolio is first to focus on the assets (equities, stocks) in the index. Then to consider for these assets the known prices (and returns) in the immediate past, these being realisations of some underlying stochastic processes. Finally this past data is used in conjunction with some optimisation model to decide a portfolio that aims to outperform the index.
@@
An implicit assumption here is that any asset considered for inclusion in the enhanced indexation portfolio based on the immediate past will also exist into the immediate future. For assets in major indices such as the S&P 500 this assumption is valid. However this assumption is no longer true if we consider options.
@@
Briefly here, since we discuss options in greater detail below, financial options exist relating to the market index value. For example a call option relating to the S&P 500 would have a specified index value ($I$, say), at a specified time in the future - so the option has a specified life before expiry. If we choose to pay to buy this option now then we can, if we wish, sell it over its life. At the end of its life, assuming we have not sold the option in the meantime, if the market index is above $I$ then we receive a cash amount depending upon the difference between the market index value and $I$. If the market index value is below $I$ then we receive nothing.
@@
But because an individual option has a specified life, and indeed can be sold over its life, the assumption referred to above, namely any asset considered for inclusion in the enhanced indexation portfolio based on the immediate past (the in-sample period) will also exist into the immediate future (the out-of-sample period), is no longer true. This is especially so for the two S&P 500 options considered in this paper which are: Monthly, expiring on the third Friday of every month and EOM, expiring on the last business day of each month. So, for example, if the in-sample period exceeds a month any option bought at the start of the period will have expired and we have the issue of what do we do with any cash generated?
@@
The difficulty here is that we have a mismatch: equity assets have a “long life”, individual options have a “short life”. In this paper we propose an approach based on option strategies that enables us to overcome this mismatch, and hence consider both equity assets and options for inclusion in an enhanced indexation portfolio.
The simple justification as to why we might consider index options in enhanced indexation is that index options on major indices are widely available. As such including them in a portfolio seeking to outperform an index can, depending upon the decisions taken, contribute to achieving better outperformance than using equity assets alone.
@@
An option strategy for a given option is a specified set of rules which detail how the option is to be traded (bought, rolled over, sold) depending upon market conditions. Applying an option strategy to the immediate past means that we can evaluate it, so effectively it is equivalent to an asset with known performance. Because it is a strategy it has a “long life”, so can be chosen for inclusion in the enhanced indexation portfolio and applied in the future. Each option strategy contributes one asset. In other words we have converted a “short life” option into a “long life” option strategy asset.
Option strategy assets, together with equity assets, can be used in any optimisation approach for enhanced indexation. In this paper we make use of an approach based upon linear programming and second-order stochastic dominance.
@@
The structure of this paper is as follows. In Section 2 we review the relevant literature as to enhanced indexation and discuss previous work dealing with portfolio optimisation involving both equities and options. We also state what we believe to be the contribution of the paper. In Section 3 we discuss the varying options that are available and how they are valued (priced), with special regard to index options for the S&P 500. We define what we mean by an option strategy, giving an example S&P 500 option strategy and illustrate how we can evaluate such a strategy.
@@ 2 Literature review
This paper deals with enhanced indexation making use of both equities and options. As far as we are aware there are no previous papers in the literature dealing with using options together with equities for enhanced indexation. Our paper is hence the first to deal with this problem. There are some papers dealing with using options together with equities for other portfolio optimisation problems and these are reviewed below.
2.1 Enhanced indexation
@@
In this section we review the literature relating to enhanced indexation discussing papers that, in our view, are of more significance. We would mention here the recent work of Silva and de Almeida Filho (2024) who gave a comprehensive literature review which includes approaches for index tracking extended to deal with enhanced indexation.
@@
Early work dealing with enhanced indexation was given by Alexander and Dimitriu (2005) who made use of cointegration and Dose and Cincotti (2005) who clustered stocks based on a distance measure between two stocks defined using the difference between their stock price histories (amongst other approaches).
@@
Canakgoz and Beasley (2009) presented a mixed-integer programming formulation that includes transaction costs, a cardinality constraint limiting the number of stocks that can be purchased and a transaction cost constraint limiting the total transaction cost that can be incurred. Their approach makes use of the linear regression between portfolio returns and index returns. Li et al. (2011) proposed a bi-objective formulation, minimising a tracking error function related to returns below the index, whilst maximising portfolio return and limiting transaction costs. Meade and Beasley (2011) presented a paper that makes use of a modified Sortino ratio to produce portfolios of specified cardinality, but with an upper limit on the holding in any particular asset.
@@
Guastaroba and Speranza (2012) presented a mixed-integer programming formulation that includes transaction costs, a cardinality constraint and a transaction cost constraint. Their objective is to minimise the absolute deviation between a scaled increased index value and the portfolio value. Mezali and Beasley (2013) presented a mixed-integer programming formulation that includes transaction costs, a cardinality constraint and a transaction cost constraint. Their objective is based on quantile regression.
@@
Roman et al. (2013) made use of second-order stochastic dominance (SSD), aiming for a portfolio that is second-order stochastic dominant with regard to the index. Their solution procedure involves the use of cutting planes as given by Fábián et al. (2011a, b) who introduced a cutting plane reformulation of Roman et al. (2006). Valle et al. (2014) presented a three-stage mixed-integer programming formulation that includes transaction costs, a cardinality constraint and a transaction cost constraint. The first two stages relate to a regression of portfolio return against time, whilst the third stage relates to minimising transaction cost.
@@
Acosta-Gonzalez et al. (2015) presented an approach based on optimising cointegration between the portfolio chosen and the benchmark index. They made use of genetic algorithms. For enhanced index tracking they artificially increased the benchmark index. Bruni et al. (2015) proposed a linear bi-objective optimisation approach that maximises average excess return and minimises the maximum underperformance with respect to the index. Filippi et al. (2016) presented a bi-objective mixed-integer programming formulation balancing excess return against the absolute deviation between a scaled index value and the portfolio value.
@@
Clark et al. (2019) presented an approach based upon reweighting the assets in the benchmark index. Their approach makes use of the concept of cumulative utility area ratio, which measures the incremental utility of one asset compared to another. Guastaroba et al. (2020) proposed bi-criteria optimisation models using a weighted combination of multiple conditional value-at-risk (CVaR) measures. Li et al. (2021) presented a nonlinear enhanced indexation model with an explicit objective to track and outperform the naive diversification strategy.
Xu et al. (2022) make use of group-sparse optimisation, with dynamic parameter adjustment being based on either a hidden Markov model or an artificial neural network. Allen-Zhao et al. (2024) presented an approach based on robust CVaR and group-sparse optimisation. Dai and Li (2024) make use of a novel neural network together with a hidden Markov model with Gaussian noise to capture regime switching. Cesarone and Puerto (2025) presented an approach based upon ordered weighted average in conjunction with stochastic dominance. The weighted average is based on the sum of $k$ maximal CVaRs.
@@
2.2 Portfolio optimisation with equities and options
In terms of papers dealing with using options together with equities for other portfolio optimisation problems a common assumption is that an option is held (unchanged) from portfolio creation until the end of the investment period. So options are only traded at portfolio creation/rebalancing, no option trading is done over the investment period - in sharp contrast to the approach put forward in this paper. In addition some papers also draw on the framework established by Markowitz (1952), balancing return against some risk measure.
Note that we only consider here work involving some optimisation approach to portfolio construction, excluding work such as Merton et al. (1978, 1982); Bookstaber and Clarke (1984) that did not optimise portfolio composition.
Liang et al. (2008) used a Markowitz framework with variance as the risk measure and presented approaches based on stochastic programming and stochastic control with dynamic programming. Scheuenstuhl and Zagst (2008) used a Markowitz framework with variance as the risk measure and introduced shortfall constraints based upon lower partial moments. They also introduced monotony constraints to prevent speculation against the general market trend.
Zymler et al. (2011) presented a robust optimisation model to maximise the worst-case portfolio return if equity returns lie within their uncertainty set. Ashrafi and Thiele (2021) presented a similar approach, but with a different uncertainty set.
Davari-Ardakani et al. (2016) used a multistage scenario tree with decisions (as to equities and options) made at each node of the tree. Their objective is to maximise expected terminal wealth (after discounting) minus the present value of expected downside deviation (weighted at each scenario node) from a target wealth level. Maasar et al. (2016, 2022) used a Markowitz framework with a target return, but varied the risk measure considering both variance and CVaR.
Zhao and Palomar (2018) extended the Markowitz framework, adopting a weighted objective balancing expected return against risk (variance) and transaction cost. Khodamoradi et al. (2020) used mean-absolute deviation together with a cardinality constraint on the number of equities held to find a portfolio involving both equities and options. He and Kwon (2023) used the Markowitz framework adopting a weighted objective balancing expected return against either variance or CVaR.
We would stress here that none of the papers reviewed in this section use options in a portfolio optimisation setting in as general a way as we propose in this paper, allowing trading of options before expiry in the investment period.
2.3 Contribution
Above we introduced the concept of an option strategy, defining for a given option a trading strategy which turns it from a “short life” option into a “long life” asset. This enables us to include options as assets in enhanced indexation. As stated previously, as far as we are aware there are no previous papers in the literature dealing with using options together with equities for enhanced indexation. Our paper is hence the first to deal with this problem.
But we believe that the contribution of this paper reaches far beyond enhanced indexation because of its potential impact. The work presented in this paper shows how options can be systematically and mathematically incorporated as assets into any portfolio optimisation problem. In the light of this we believe that the contribution of this paper is:
  • to introduce the concept of an option strategy
  • to, for the first time in the literature, include options as assets in enhanced indexation
  • to give computational results showing that introducing option strategies in an enhanced indexation setting offers clear benefits in terms of improved out-of-sample performance. This applies whether we use equities or an exchange-traded fund as part of the enhanced indexation portfolio.
  • to enable options (not just index options) to be included in many other situations. For example:
    • making use of call and put options for individual equity assets, e.g. those that are traded on major companies in an index
    • considering other types of options, e.g. futures
    • introducing options into any portfolio optimisation problem that involves equity assets, e.g. Markowitz based approaches balancing return against some risk measure
    • trading options over time in an attempt to make money, but without involving equity assets
@@ 3 Options
In this section we discuss the varying options that are available and how they are valued (priced), with special regard to index options for the S&P 500, which we use for our computational work below. We then define what we mean by an option strategy, giving an example S&P 500 option strategy and illustrate how we can evaluate such a strategy. We go on to discuss how we can use option strategies in any optimisation based approach to enhanced indexation. Finally we discuss the choice of an optimisation based enhanced indexation approach.
@@ 3.1 Options and option pricing
A call option on an underlying asset gives the holder of the option the right (but not the obligation) to buy the asset at a fixed price, called the strike (or exercise) price. If this right can only be exercised at a specific (expiry) time in the future then it is known as an European call option. If it can be exercised at any time up to expiry then it is known as an American call option.
A put option is the right to sell the asset, otherwise it has the same characteristics as a call option (so a strike price, expiry time, European or American).
@@
Even if an option can only be exercised at expiry it can be bought and sold at any time. So for example if we buy a call or put option now we can sell it at any time between buying it and expiry.
@@
However options also apply if we have intangible assets, such as the S&P 500 index. As we cannot “buy” any of the index an option such as this is cash settled. So the strike price relates to the level of the index. Considering a call option, if at expiry the index is below the strike price the option is worthless. However if at expiry the index is above the strike price then the option is cash settled, so the option holder receives a cash amount linearly related to the difference between the current index level and the strike price.
@@ 3.2 Defining an option strategy
Collectively these assumptions mean that in any time period we are either holding cash or the option. On a technical note here although we use phrases such as “buy the option” we mean buy a number of units of the same option. Since the sums of money involved are likely to be large we can ignore any issues relating to fractional option unit purchase.
@@
In other words an option strategy is equivalent to an asset that can be treated just like an equity asset in any current approach for enhanced indexation.
@@ 3.3 An example S&P 500 option strategy
  • In terms of expiry dates the two most liquid S&P 500 options are: Monthly, expiring on the third Friday of every month and EOM, expiring on the last business day of each month.
@@
To illustrate how we rollover suppose the rollover rule above is active. So we are selling the current put option and buying another put option. For the purchased option we need to decide the lifetime and the exercise price. As the rollover rule involves $L<20$ days to invoke sale of the current option then for the purchased option we choose the next available expiry (monthly or EOM) 20 or more days in the future, giving the purchased option a lifetime of expiry time minus current time.
In terms of the exercise price $E$ for the purchased OTM option then as we have a $3\%$ moneyness target and since for a OTM put option $E<F_{t}$ (see Table 1) we have $100(F_{t}-E)/F_{t}=3$ from Equation (6), equivalently $E=0.97F_{t}$. Here $F_{t}$ is calculated from Equation (5) using the lifetime as decided above for the purchased option. As exercise prices are integer multiples of 5 we round $E$ to the nearest multiple of 5.
@@ 3.4 Example S&P 500 option strategy evaluation
In period 2 the 30 day index return is $\leq-5\%$ so we buy the option at 90 so in total $1000u/90$ options are bought, with 21 days to expiry.
@@
In period 4 the options held in period 3 are now 19 days to expiry so we rollover, selling our current holding of $1000u/90$ options for 80. For the new option the number of days to expiry is 29 (e.g. EOM). Since, based on this new lifetime of 29 days, $F_{t}=6379$ the exercise price $E=0.97F_{t}$, which after after rounding to the nearest multiple of 5 is 6190. The new option costs 75 and we hold $(80{\times}1000u)/(90{\times}75)$ units of that option.